OpenStudy (anonymous):

A copy machine makes 28 copies per minute. How long does it take to make 133 copies?

OpenStudy (anonymous):

??

OpenStudy (unklerhaukus):

the rate of copies is \[r\left[\frac{\text{copies}}{\text{minute}}\right]=\frac {C[\text{copies}]}{t[{\text{min}}]}\] \[\text{rate of copy machine}\quad r\left[\frac{\text{copies}}{\text{minute}}\right]=\quad28\left[\frac{\text{copies}}{\text{minute}}\right]\]

OpenStudy (unklerhaukus):

\[r\left[\frac{\text{copies}}{\text{minute}}\right]=\frac {C[\text{copies}]}{t[{\text{min}}]}\] \[r=\frac Ct\] \[t=\frac{C}{r}\]

OpenStudy (unklerhaukus):

\[t[{\text{min}}]=\frac {C[\text{copies}]}{r\left[\frac{\text{copies}}{\text{minute}}\right]}=\frac {133[\text{copies}]}{28\left[\frac{\text{copies}}{\text{minute}}\right]}=\]